# A new (but very nearly old) proof of the Rogers–Ramanujan identities

### v2

@@ 3. Some additional results
$f(z)=(-z;q)_{\infty}=\sum_{k=0}^{\infty}\frac{q^{k(k-1)/2}z^{k}}{(q;q)_{k}}$
@@
$\theta(1/z;q)(qtz;q)_{\infty}$ $=\frac{\theta(1/z;q)}{(-qtz;q)_{\infty}}\,(q^{2}t^{2}z^{2};q^{2})_{\infty}$
$=\frac{\theta(1/z;q)}{(q^{1/2}tz;q)_{\infty}}\,(q^{1/2}tz;q^{1/2})_{\infty}=\frac{\theta(1/z;q)}{(q^{3/2}tz;q)_{\infty}}\,(qtz;q^{1/2})_{\infty}$
@@
$R(t;q)$ $=(-qt;q)_{\infty}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{3k^{2}}t^{2k}}{(q^{2};q^{2})_{k}(-qt;q)_{2k}}$ (3.2a)
$=(q^{1/2}t;q)_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k-1)/4}t^{k}}{(q^{1/2};q^{1/2})_{k}(q^{1/2}t;q)_{k}}$ (3.2b)
$=(q^{3/2}t;q)_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k+1)/4}t^{k}}{(q^{1/2};q^{1/2})_{k}(q^{3/2}t;q)_{k}}.$ (3.2c)
The identity (3.2a) is the limit case $B\rightarrow\infty$ of [GS, (7.9)] and (3.2b) is [GS, (7.17)]. We have not found (3.2c) in the literature but it can be deduced from (3.2b) by a simple contiguity argument. Namely, splitting the terms in (3.2c) using
$(q^{3/2}t;q)_{\infty}\frac{t^{k}}{(q^{3/2}t;q)_{k}}=(q^{1/2}t;q)_{\infty}\left(\frac{t^{k}}{(q^{1/2}t;q)_{k}}+\frac{q^{k+\frac{1}{2}}t^{k+1}}{(q^{1/2}t;q)_{k+1}}\right)$
and then replacing $k$ by $k-1$ in the second sum one finds that (3.2c) equals (3.2b).
In the cases $t=1$ and $t=q$, (3) can be written as
@@
$=(q^{1/2};q)_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k-1)/4}}{(q^{1/2};q^{1/2})_{k}(q^{1/2};q)_{k}},$ (3.3b)
$=(q^{1/2};q)_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k+1)/4}}{(q^{1/2};q^{1/2})_{k}(q^{1/2};q)_{k+1}}$ (3.3c)
@@
$=(q^{1/2};q)_{\infty}\sum_{k=0}^{\infty}\frac{q^{3k(k+1)/4}}{(q^{1/2};q^{1/2})_{k}(q^{1/2};q)_{k+1}}$ (3.3e)
$=(q^{1/2};q)_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k+5)/4}}{(q^{1/2};q^{1/2})_{k}(q^{1/2};q)_{k+2}}.$ (3.3f)
The summations (3.3a)–(3.3e) are, respectively, 19, 46, 62, 97 and 44 on Slater’s list (19 is equivalent to 100, and 44 to 63). Most of them, (3.3a)–(3.3c) and (3.3e), were obtained by Rogers [R1, R3], whereas the earliest reference for (3.3d) is Ramanujan’s lost notebook [AB, (11.2.7)]. We have not found (3.3f) in the literature, but as explained above it follows easily from (3.3e).
@@
$H(q)=(-q;q)_{\infty}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{k(3k-2)}}{(q^{4};q^{4})_{k}(-q;q^{2})_{k}},$ (3.3g)
see [AB, p. 253] for the derivation of (3.3d) from (3.3g). Not surprisingly, one can obtain (3.3g) by a slight modification of our proof of (3.3d). We provide some details since it is another nice example of Andrews’ constant term method. We first write
$\frac{(-1)^{k}q^{3k(k-2)}}{(q^{4};q^{4})_{k}(-q;q^{2})_{k}}=\frac{(-1)^{k}q^{k(k-1)}}{(q^{2};q^{2})_{k}}\cdot\frac{(-1)^{2k}q^{2k(2k-1)/2}}{(-q;q)_{2k}},$
@@
$(-q;q)_{\infty}\operatorname{CT}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{k(k-1)}z^{2k}}{(q^{2};q^{2})_{k}}\sum_{l=-\infty}^{\infty}\frac{(-1)^{l}q^{l(l-1)/2}}{(-q;q)_{l}z^{l}}\\ =\operatorname{CT}(z^{2};q^{2})_{\infty}\frac{\theta(1/z;q)}{(-qz;q)_{\infty}}=\operatorname{CT}(1+z)(z;q)_{\infty}\theta(1/z;q).$
@@
$\operatorname{CT}(1+z)(z;q)_{\infty}\theta(1/z;q)=\operatorname{CT}(z;q)_{\infty}(\theta(1/z;q)-\theta(q/z;q))\\ =\operatorname{CT}((z;q)_{\infty}-(qz;q)_{\infty})\theta(1/z;q)=R(q^{-1};q)-R(1;q).$
@@
$\sum_{k=0}^{\infty}\frac{q^{k^{2}}(q^{-k}-1)}{(q;q)_{k}}=\sum_{k=1}^{\infty}\frac{q^{k^{2}-k}}{(q;q)_{k-1}}=\sum_{k=0}^{\infty}\frac{q^{k^{2}+k}}{(q;q)_{k}}$
@@
Constant-term proofs of the identities (3.3a) and (3.3g) were previously given by Sills [S1]. He uses the series (2.8) rather than the original Rogers–Ramanujan identities (1).
4. Rogers’ proof of the quadratic transformations
Since Rogers’ papers are not easy to read, we will provide a slightly modified version of his proof of the quadratic transformations (2). Although the notation was not used by Rogers, we find it instructive to introduce the $q$-exponential functions
$e_{q}(x)$ $=\frac{1}{(x;q)_{\infty}}=\sum_{k=0}^{\infty}\frac{x^{k}}{(q;q)_{k}},$
$E_{q}(x)$ $=(-x;q)_{\infty}=\sum_{k=0}^{\infty}\frac{q^{k(k-1)/2}x^{k}}{(q;q)_{k}}.$
@@ 4. Rogers’ proof of the quadratic transformations
It is easy to see that
$\delta_{x}e_{q}(tx)=te_{q}(tx)$ (4.1)
and hence
$f(\delta_{x})e_{q}(tx)=f(t)e_{q}(tx)$
for any formal power series $f$. In particular,
@@
In [R1], Rogers proves (4.3) in a more direct way and uses it to derive (4.2).
The second $q$-exponential function satisfies
$\delta_{x}E_{q}(tx)=tE_{q}(qtx),$
which gives
$\delta_{x}^{k}E_{q}(tx)=q^{k(k-1)/2}t^{k}E_{q}(q^{k}tx)$
and
@@
Specializing $t=q$ we obtain [R2, §6, Eq. (1)] in the form
@@
$\sum_{k=0}^{\infty}H_{k}(x,xq^{1/2})t^{k}=\frac{1}{(tx;q^{1/2})_{\infty}},$
@@
$H_{k}(x,xq^{1/2})=\frac{x^{k}}{(q^{1/2};q^{1/2})_{k}}.$
@@
$\sum_{k=0}^{\infty}\frac{q^{k(k+1)}x^{k}}{(q;q)_{k}}=(-q^{2}x;q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+2)}x^{k}}{(-q^{2}x,q^{2};q^{2})_{k}}.$ (4.6a)
@@
$\sum_{k=0}^{\infty}\frac{q^{k(k+1)}x^{k}}{(q;q)_{k}}=(-q^{3}x;q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}x^{k}}{(-q^{3}x,q^{2};q^{2})_{k}}.$ (4.6b)
@@
$\sum_{k=0}^{\infty}H_{k}(x,-x)t^{k}=\frac{1}{(t^{2}x^{2};q^{2})_{\infty}},$
@@
$H_{2k}(x,-x)=\frac{x^{2k}}{(q^{2};q^{2})_{k}},\qquad H_{2k+1}(x)=0.$
@@
$\sum_{k=0}^{\infty}\frac{q^{2k(2k+1)}x^{2k}}{(q^{2};q^{2})_{k}}=(-qx;q)_{\infty}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{k(k+1)/2}x^{k}}{(-qx,q;q)_{k}}.$ (4.6c)
The identities (4) are indeed equivalent to (2).

### v3

@@ 3 Some additional results
$f(z)=(-z;q)_{\infty}=\sum_{k=0}^{\infty}\frac{q^{k(k-1)/2}z^{k}}{(q;q)_{k}},$
@@
$\theta(1/z;q)(qtz;q)_{\infty}$ $=\frac{\theta(1/z;q)}{(-qtz;q)_{\infty}}\bigl{(}q^{2}t^{2}z^{2};q^{2}\bigr{)}_{\infty}$
$=\frac{\theta(1/z;q)}{\bigl{(}q^{1/2}tz;q\bigr{)}_{\infty}}\bigl{(}q^{1/2}tz;q^{1/2}\bigr{)}_{\infty}=\frac{\theta(1/z;q)}{\bigl{(}q^{3/2}tz;q\bigr{)}_{\infty}}\bigl{(}qtz;q^{1/2}\bigr{)}_{\infty}$
@@
$R(t;q)$ $=(-qt;q)_{\infty}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{3k^{2}}t^{2k}}{\bigl{(}q^{2};q^{2}\bigr{)}_{k}(-qt;q)_{2k}}$ (3.2a)
$=\bigl{(}q^{1/2}t;q\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k-1)/4}t^{k}}{\bigl{(}q^{1/2};q^{1/2}\bigr{)}_{k}\bigl{(}q^{1/2}t;q\bigr{)}_{k}}$ (3.2b)
$=\bigl{(}q^{3/2}t;q\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k+1)/4}t^{k}}{\bigl{(}q^{1/2};q^{1/2}\bigr{)}_{k}\bigl{(}q^{3/2}t;q\bigr{)}_{k}}.$ (3.2c)
The identity (3.2a) is the limit case $B\rightarrow\infty$ of [9, equation (7.9)] and (3.2b) is [9, equation (7.17)]. We have not found (3.2c) in the literature but it can be deduced from (3.2b) by a simple contiguity argument. Namely, splitting the terms in (3.2c) using
$\bigl{(}q^{3/2}t;q\bigr{)}_{\infty}\frac{t^{k}}{\bigl{(}q^{3/2}t;q\bigr{)}_{k}}=\bigl{(}q^{1/2}t;q\bigr{)}_{\infty}\left(\frac{t^{k}}{\bigl{(}q^{1/2}t;q\bigr{)}_{k}}+\frac{q^{k+\frac{1}{2}}t^{k+1}}{\bigl{(}q^{1/2}t;q\bigr{)}_{k+1}}\right)$
and then replacing $k$ by $k-1$ in the second sum, one finds that (3.2c) equals (3.2b).
In the cases $t=1$ and $t=q$, (3.2) can be written as
@@
$=\bigl{(}q^{1/2};q\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k-1)/4}}{\bigl{(}q^{1/2};q^{1/2}\bigr{)}_{k}\bigl{(}q^{1/2};q\bigr{)}_{k}}$ (3.3b)
$=\bigl{(}q^{1/2};q\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k+1)/4}}{\bigl{(}q^{1/2};q^{1/2}\bigr{)}_{k}\bigl{(}q^{1/2};q\bigr{)}_{k+1}}$ (3.3c)
@@
$=\bigl{(}q^{1/2};q\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{3k(k+1)/4}}{\bigl{(}q^{1/2};q^{1/2}\bigr{)}_{k}\bigl{(}q^{1/2};q\bigr{)}_{k+1}}$ (3.3e)
$=\bigl{(}q^{1/2};q\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(3k+5)/4}}{\bigl{(}q^{1/2};q^{1/2}\bigr{)}_{k}\bigl{(}q^{1/2};q\bigr{)}_{k+2}}.$ (3.3f)
The summations (3.3a)–(3.3e) are, respectively, 19, 46, 62, 97 and 44 on Slater’s list (19 is equivalent to 100, and 44 to 63). Most of them, (3.3a)–(3.3c) and (3.3e), were obtained by Rogers [14, 16], whereas the earliest reference for (3.3d) is Ramanujan’s lost notebook [4, equation (11.2.7)]. We have not found (3.3f) in the literature, but as explained above it follows easily from (3.3e).
@@
$H(q)=(-q;q)_{\infty}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{k(3k-2)}}{\bigl{(}q^{4};q^{4}\bigr{)}_{k}\bigl{(}-q;q^{2}\bigr{)}_{k}},$ (3.3g)
see [4, p. 253] for the derivation of (3.3d) from (3.3g). Not surprisingly, one can obtain (3.3g) by a slight modification of our proof of (3.3d). We provide some details since it is another nice example of Andrews’ constant term method. We first write
$\frac{(-1)^{k}q^{3k(k-2)}}{\bigl{(}q^{4};q^{4}\bigr{)}_{k}\bigl{(}-q;q^{2}\bigr{)}_{k}}=\frac{(-1)^{k}q^{k(k-1)}}{\bigl{(}q^{2};q^{2}\bigr{)}_{k}}\cdot\frac{(-1)^{2k}q^{2k(2k-1)/2}}{(-q;q)_{2k}},$
@@
$(-q;q)_{\infty}\operatorname{CT}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{k(k-1)}z^{2k}}{\bigl{(}q^{2};q^{2}\bigr{)}_{k}}\sum_{l=-\infty}^{\infty}\frac{(-1)^{l}q^{l(l-1)/2}}{(-q;q)_{l}z^{l}}$
$\qquad=\operatorname{CT}\bigl{(}z^{2};q^{2}\bigr{)}_{\infty}\frac{\theta(1/z;q)}{(-qz;q)_{\infty}}=\operatorname{CT}(1+z)(z;q)_{\infty}\theta(1/z;q).$
@@
$\operatorname{CT}(1+z)(z;q)_{\infty}\theta(1/z;q)$ $=\operatorname{CT}(z;q)_{\infty}(\theta(1/z;q)-\theta(q/z;q))$
$=\operatorname{CT}((z;q)_{\infty}-(qz;q)_{\infty})\theta(1/z;q)=R\bigl{(}q^{-1};q\bigr{)}-R(1;q).$
@@
$\sum_{k=0}^{\infty}\frac{q^{k^{2}}\bigl{(}q^{-k}-1\bigr{)}}{(q;q)_{k}}=\sum_{k=1}^{\infty}\frac{q^{k^{2}-k}}{(q;q)_{k-1}}=\sum_{k=0}^{\infty}\frac{q^{k^{2}+k}}{(q;q)_{k}},$
@@
Constant-term proofs of the identities (3.3a) and (3.3g) were previously given by Sills [18]. He uses the series (2.7) rather than the original Rogers–Ramanujan identities (1.1).
4 Rogers’ proof of the quadratic transformations
Since Rogers’ papers are not easy to read, we will provide a slightly modified version of his proof of the quadratic transformations (2.5). Although the notation was not used by Rogers, we find it instructive to introduce the $q$-exponential functions
$e_{q}(x)$ $=\frac{1}{(x;q)_{\infty}}=\sum_{k=0}^{\infty}\frac{x^{k}}{(q;q)_{k}},\qquad E_{q}(x)=(-x;q)_{\infty}=\sum_{k=0}^{\infty}\frac{q^{k(k-1)/2}x^{k}}{(q;q)_{k}}.$
@@ 4 Rogers’ proof of the quadratic transformations
It is easy to see that $\delta_{x}e_{q}(tx)=te_{q}(tx)$, and hence $f(\delta_{x})e_{q}(tx)=f(t)e_{q}(tx)$ for any formal power series $f$. In particular,
@@
In [14], Rogers proves (4.2) in a more direct way and uses it to derive (4.1).
The second $q$-exponential function satisfies $\delta_{x}E_{q}(tx)=tE_{q}(qtx)$, which gives $\delta_{x}^{k}E_{q}(tx)=q^{k(k-1)/2}t^{k}E_{q}\bigl{(}q^{k}tx\bigr{)}$ and
@@
Specializing $t=q$, we obtain [15, Section 6, equation (1)] in the form
@@
$\sum_{k=0}^{\infty}H_{k}\bigl{(}x,xq^{1/2}\bigr{)}t^{k}=\frac{1}{\bigl{(}tx;q^{1/2}\bigr{)}_{\infty}},$
@@
$H_{k}\bigl{(}x,xq^{1/2}\bigr{)}=\frac{x^{k}}{\bigl{(}q^{1/2};q^{1/2}\bigr{)}_{k}}.$
@@
$\sum_{k=0}^{\infty}\frac{q^{k(k+1)}x^{k}}{(q;q)_{k}}=\bigl{(}-q^{2}x;q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+2)}x^{k}}{\bigl{(}-q^{2}x,q^{2};q^{2}\bigr{)}_{k}}.$ (4.5a)
@@
$\sum_{k=0}^{\infty}\frac{q^{k(k+1)}x^{k}}{(q;q)_{k}}=\bigl{(}-q^{3}x;q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}x^{k}}{\bigl{(}-q^{3}x,q^{2};q^{2}\bigr{)}_{k}}.$ (4.5b)
@@
$\sum_{k=0}^{\infty}H_{k}(x,-x)t^{k}=\frac{1}{\bigl{(}t^{2}x^{2};q^{2}\bigr{)}_{\infty}},$
@@
$H_{2k}(x,-x)=\frac{x^{2k}}{\bigl{(}q^{2};q^{2}\bigr{)}_{k}},\qquad H_{2k+1}(x,x)=0.$
@@
$\sum_{k=0}^{\infty}\frac{q^{2k(2k+1)}x^{2k}}{\bigl{(}q^{2};q^{2}\bigr{)}_{k}}=(-qx;q)_{\infty}\sum_{k=0}^{\infty}\frac{(-1)^{k}q^{k(k+1)/2}x^{k}}{(-qx,q;q)_{k}}.$ (4.5c)
The identities (4.5) are indeed equivalent to (2.5).
Acknowledgements
It is a pleasure to dedicate this paper to Stephen Milne, whose work has been one of my main sources of inspiration. I would like to thank Andrew Sills for the reference to his master’s thesis and for making it available [18]. I also thank the referees for several useful comments. This research is supported by the Swedish Science Research Council, project no. 2020-04221.